OpenStudy (anonymous):

What is the percent decrease from 60 to 39?

OpenStudy (igreen):

First subtract them, 60 - 39 = ?

OpenStudy (anonymous):

I think it Is 35

OpenStudy (anonymous):

I did what we did before

OpenStudy (anonymous):

Subtract then divide then multiply

OpenStudy (igreen):

Correct.

OpenStudy (igreen):

35 / 60 = 0.35 0.35 * 100 = 35

OpenStudy (anonymous):

Yay! You taught me how to figure these things out!